Back to main Archive News page.
Mon 31 Jul 2006 - All systems ran without problems
Fri 28 Jul 2006 - New Exchange Email Systems
Following the successful deployment of new Exchange Email hardware, ISS will
be migrating all current Exchange users to the newer systems. This process may
cause a prompt to appear, which states:
"The Microsoft Exchange administrator has made a change that requires you quit and restart Outlook". If you see this prompt, please restart Outlook, as requested.
This process will not cause the loss of any email. Any other problems may be referred to the ISS Service Desk (x10987).
Thu 27 Jul 2006 - Core Switch Problems
Due to continuing problems with the core switch in the Computer Centre it was necessary to replace part of the hardware. Work was scheduled to commence at 18:00 and last for approximately 45 minutes, however, work commenced 20 minutes early due to the untimely failure of the equipment. ISS continue to monitor the situation for further problems.
Wed 26 Jul 2006 - All systems ran without major problems
Tue 25 Jul 2006 - Network problems, Network Maintenance & Fraudulent email
At approx 23:00hrs one of the 3 core switches failed causing an outage to a number of areas across campus. ISS Network engineers restored full service by 23:20.
A convincing fake mail message appearing to come from PayPal, with a subject line containing words like "Unauthorized use of your PayPal account" is circulating. The message does not come from PayPal; and as with all such "phishing" messages, you should never enter details of accounts when asked.
08:00-09:00 Two network hardware upgrades and replacements took place. The first was in the computer centre. The second replacement was a complete hardware upgrade for some equipment feeding wide areas of north campus residences which has caused short periods of network problems over the last week.
Mon 24 Jul 2006 - All systems ran without major problems
Fri 21 Jul 2006 - All systems ran without major problems
Thu 20 Jul 2006 - All systems ran without major problems
Wed 19 Jul 2006 - All systems ran without major problems
14:00-15:00 LUVLE and My Modules was unavailable from 2pm on Wednesday 19th July for approximately one hour, to allow for essential server maintenance work to be carried out.
Tue 18 Jul 2006 - Loss of network connectivity on campus
15:00 Network connectivity in some parts of campus was lost for about 10 minutes today, 18/7/06 at about 3pm. This was due to a core switch failure.
Mon 17 Jul 2006 - All systems ran without major problems
Fri 14 Jul 2006 - All systems ran without major problems
A number of blocks in Graduate College experienced connectivity problems. Update 09:29 - Networking have resolved the problems seen in Graduate although the cause of the problems were not clear. Please contact the ISS Service Desk x10987, if any more problems are seen
Thu 13 Jul 2006 - All systems ran without major problems
Wed 12 Jul 2006 - Central Filestore Maintenance
13:00-14:30 The central filestore (H: drive) and departmental filestore were unavailable due to routine maintenance. All other systems ran without problems
Tue 11 Jul 2006 - ISS Grizedale PC Lab A240 Closure
The ISS open access PC Lab in Grizedale A240 has been decommissioned. Alternative accommodation at the south end of campus for open access PCs is being sought for the start of the 2006/07 academic year. See locations of other Labs on campus.
Mon 10 Jul 2006 - All systems ran without major problems
Fri 07 Jul 2006 - All systems ran without major problems
Thu 06 Jul 2006 - All systems ran without major problems
Wed 05 Jul 2006 - Minor network problems in the Library
Users in the Library experienced intermittent network connectivity after a failure with a network switch. The problem was quickly isolated and full connectivity restored within a few minutes.
Tue 04 Jul 2006 - Network Maintenance and Login problems
During the 'Network at Risk' period earlier today, UKERNA performed a number of tests on the Internet connection to the University.
There was a problem with the DHCP server which meant that users would be unable to login. The problem with DHCP was resolved, but some users found they had to reboot their machine in order to get a valid DHCP lease. Users experiencing problems should contact the ISS Staff Help Desk on x10987